i. The one-year interest rate in the US. is is-200 and in the euro zone the one-year interest rate is ie = 6%. The one-year forward exchange rate is $1.20 = €1.00, what must the spot rate be to eliminate arbitrage opportunities? A), є 1.2471-$1.00 B). € 1.00 = $ 1.2471 C).1.00 $1.1547
Answer #1

US Interest Rate = i(US) = 2 % and Euro Interest Rate = i(EU) = 6 %, One - Year Forward Rate = $ 1.2 / EUR

Let the current spot rate be $ K / EUR

Then, as per interest rate parity 1.2 = [1+i(US) / 1+i(EU)] x K = [1.02/1.06] x K

K = [1.06/1.02] x 1.2 = 1.247059 ~ 1.2471 $ / EUR

Therefore, the correct option is (B)